Thread View: sci.space.news
1 messages
1 total messages
Started by Ron Baalke
Thu, 21 Jan 1993 18:53
Ulysses Update - 01/18/93
Author: Ron Baalke
Date: Thu, 21 Jan 1993 18:53
Date: Thu, 21 Jan 1993 18:53
125 lines
4716 bytes
4716 bytes
Forwarded from Peter Beech, Ulysses Mission Operations Manager
SUBJECT: ULYSSES MISSION OPERATIONS REPORT No. 125
PERIOD: 12th January to 18th January 1993
1. MISSION OPERATIONS
Experiment reconfigurations have been carried out as
required.
A HUS (Solar X-ray instrument) calibration test was conducted on
12th January.
A HED (Magnetic Field instrument) calibration was conducted on
13th January.
A KEP EPAC (Energetic Particles instrument) in-flight calibration
was conducted on 18th January.
Routine Earth pointing manoeuvres were performed on 14th and
17th January.
An average of 92.06 % data recovery was achieved during the
reporting period.
The percentage of data acquired as a function of bit rate is
as follows:
1024 bps 49.63 %
512 bps 42.37 %
The lower than normal data recovery for this reporting
period was due to a maser low noise amplifier failure at
Goldstone which caused a complete loss of a scheduled pass.
The outage was minimised by support being provided from the
Canberra station. A total of 10 hours of real time data
were lost.
2. SPACECRAFT STATUS
POWER
Nominal.
Estimated S/C power consumption 257 watts.
AOCS
Nominal.
TTC
The spacecraft is currently configured with receiver 2 as
the prime unit fed via the high gain antenna and with
receiver 1 as backup fed through the low gain antenna (LGA-
F). The downlink is provided through EPC2/TWTA2.
The 34 meter ground stations are in use to support TTC
operations. Ranging from 34 meter ground stations is
performed when the spacecraft is configured for bit rates of
1024 bps or less. Ranging passes from 70 meter ground
stations are also being taken periodically.
Received downlink level -138 dBm.(34 meter) X-band.
(Variations of up to 10 db. can be expected as a result of
station antenna in use, local weather conditions, and
spacecraft antenna off-pointing).
Received uplink level -123 dBm.
DATA HANDLING
Nominal.
THERMAL
Nominal.
3. FLIGHT DYNAMICS
Solar Aspect The angle reduced from 9.85 on
Angle (deg.) 12th January to 9.24 on 18th January.
Sun-Probe-Earth The angle reduced from 9.81 on
Angle (deg.) 12th January to 9.12 on 18th January.
Spin Rate 4.980 rpm.
4. ORBITAL DATA
Data taken at 13:00 PDT on 18th January.
Distance from Earth 657,698,384 km.
Velocity relative to the Earth 124,022 km/hr.
Velocity relative to the Sun 34,220 km/hr.
Ecliptic latitude 16.7 deg/south
5. PLANNED OPERATIONS
Routine data gathering operations will continue together
with experiment reconfigurations as required.
On 19th January a KEP GAS (Interstellar Neutral Gas instrument)
monthly health check will be carried out.
On 25th January a KEP EPAC in-flight calibration will be
carried out.
Earth pointing manoeuvres will be performed on 20th and 23rd
January.
Periods of 2048 bps at low antenna elevations will continue
to be avoided to improve the signal to noise ratio. This
will slightly reduce the percentage of 1024 bps real-time
data received.
6. GROUND SEGMENT
The ground segment performed nominally during the reporting
period with the exception of the Goldstone maser problem
referred to above.
___ _____ ___
/_ /| /____/ \ /_ /| Ron Baalke | baalke@kelvin.jpl.nasa.gov
| | | | __ \ /| | | | Jet Propulsion Lab |
___| | | | |__) |/ | | |__ M/S 525-3684 Telos | Every once in a while,
/___| | | | ___/ | |/__ /| Pasadena, CA 91109 | try pushing your luck.
|_____|/ |_|/ |_____|/ |
Thread Navigation
This is a paginated view of messages in the thread with full content displayed inline.
Messages are displayed in chronological order, with the original post highlighted in green.
Use pagination controls to navigate through all messages in large threads.
Back to All Threads